OpenStudy (anonymous):

A line passes through (4, 5) and (8, 9). Which equation best represents the line? A. y = x + 1 B.y = 1 C.y = 2x + 1 D. y = 5x + 8

OpenStudy (ashes.boom):

do you know how to find slope

OpenStudy (anonymous):

NO

OpenStudy (ashes.boom):

slope = y2-y1/x2 -x1

OpenStudy (ashes.boom):

every coordinate has (x, y) values so the first point will be (x1, y1) and the second (x2, y2)

OpenStudy (ashes.boom):

when you plug it in what do you get?

OpenStudy (ashes.boom):

okay shes offline... slope = y2 - y1/ x2 - x1 = \[\frac{ 9-5 }{ 8-4 }\]\[\frac{ 4 }{ 4 }\]

OpenStudy (ashes.boom):

the slope is 1 so just look for the equation with a slope of 1

OpenStudy (ashes.boom):

A is the answer

OpenStudy (anonymous):

why wouldn't it be B

OpenStudy (ashes.boom):

becuase there is no slope in b. The equation of a line is y=mx+b where m is the slope and b is the y intercept. b just shows the y intercept so it would be a horizontal line with no slope

OpenStudy (ashes.boom):

does that make sense?

OpenStudy (anonymous):

yes thank you

OpenStudy (ashes.boom):

your welcome :)

OpenStudy (anonymous):

can you help me with another one

OpenStudy (ashes.boom):

sure just tag me in it :)